OpenStudy (anonymous):

how do you simplify the following. The answer is kg but i dont see how to get that. (m/s^2)(m)^2 ------------- (m^2/kg^2)/(kg-m/s^2)

OpenStudy (anonymous):

(m^2/Kg^2)/(Kg-m/s^2) =(m^2Kg-m/s^2)/Kg^2 =(m^3/s^2)/Kg (m/s^2)m^2=m^3/s^2 now (m/s^2)m^2 ---------------- (m^2/Kg^2)/(Kg-m/s^2) m^3/s^2 = ---------- (m^3/s^2)/Kg =Kg